Let T be the linear operator on R2 defined by

T(x, y) = (−y, x)

i. What is the matrix of T in the standard ordered basis for R2 ?

ii. What is the matrix of T in the ordered basis B = {α1, α2}, where α1 = (1, 2) and α2 = (1, −1)?

iii. Prove that for every real number c the operator (T − cI) is invertible.


1
Expert's answer
2021-01-08T14:39:27-0500

Here, we have

"T(x,y)=(-y,x)=\begin{bmatrix} -y\\x \end{bmatrix}"

(i) Standard ordered basis for "\> \R^2 \> is \begin{bmatrix} 1 & 0\\ 0 & 1 \end{bmatrix}"

So, for the first column applying the linear operator T we have "\begin{bmatrix} 0\\1 \end{bmatrix}" and for the second column after applying the linear operator T we have "\begin{bmatrix} -1\\0 \end{bmatrix}" . So, the matrix for T is A(let):-


"A=\begin{bmatrix}
 0 & -1 \\
 1 & 0
\end{bmatrix}"




(ii) New basis is:-


"B=\begin{bmatrix}
 1 & 1 \\
 2 & -1
\end{bmatrix}"

Now, we have to find the matrix A wrt the basis B.

Now we take the first column of B and left multiply with A.


"A\begin{bmatrix} 1\\2 \end{bmatrix}=\begin{bmatrix}
 0 & -1 \\
 1 & 0
\end{bmatrix}\cdot \begin{bmatrix} 1\\2 \end{bmatrix}=\begin{bmatrix} -2\\1 \end{bmatrix}"

Now, solving for a, b where


"\begin{bmatrix} -2\\1 \end{bmatrix}=a\begin{bmatrix} 1\\2 \end{bmatrix}+b\begin{bmatrix} 1\\-1 \end{bmatrix}\\
\Rightarrow \begin{bmatrix} -2\\1 \end{bmatrix}=\begin{bmatrix} a+b\\2a-b \end{bmatrix}"

Solving we get, a="\frac{-1}{3}" and b="\frac{-5}{3}" .


Similarly we do the above steps with the second column of B and find c, d:-


"\begin{bmatrix} 1\\1 \end{bmatrix}=\begin{bmatrix} c+d\\2c-d \end{bmatrix}"

Solving, we have c="\frac{2}{3}" and d="\frac{1}{3}" .


Therefore the final matrix wrt, the given basis B is:-


"\begin{bmatrix}
 \frac{-1}{3} & \frac{2}{3} \\\\
 \frac{-5}{3} & \frac{1}{3}
\end{bmatrix}"


(iii) Here we calculate the "|T-cI|"



"|T-cI|\\~\\
= |\begin{bmatrix}
 0 & -1 \\
 1 & 0
\end{bmatrix}-c\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}|\\~\\
=|\begin{bmatrix}
 0 & -1 \\
 1 & 0
\end{bmatrix}-\begin{bmatrix}
 c & 0 \\
 0 & c
\end{bmatrix}|\\~\\
=|\begin{bmatrix}
 -c & -1 \\
 1 & -c
\end{bmatrix}|\\~\\
=(-c)^2-(-1)\\~\\
=c^2+1\\"

Now the determinant of the matrix "T-cI" is "1+c^2" which is never equal to 0, as "c \in\R" .

Therefore the matrix "T-cI" is always invertible "\forall c\in\R" .